OpenStudy (anonymous):

4x - 3y = 17 5x + 4y = 60

OpenStudy (anonymous):

take the first equation 4x-3y=17 solve for x or y... I'll solve for X 4x=17+3y x=17/4+3/4 *y now substitute into the second equation... 5x+4y=60 5(17/4+3/4 *y)+4y=60 ... go on

OpenStudy (anonymous):

multiply both -3y and 4y with 4 and 3 respectively. Then cancel out. You'll get 16x = 68 15x = 120 add them you'll get 31x = 188 x = 188/31 then insert the value of x into one of those eqns. and you'll get the answer of y.

OpenStudy (anonymous):

sorry i meant 15x = 180 x gives 248/31 x = 8